# Refinements of the Blanco-Koldobsky-Turnšek Theorem

### v1

@@ Abstract.
We refine the well-known Blanco-Koldobsky-Turnšek Theorem which states that a norm one linear operator defined on a Banach space is an isometry if and only if it preserves orthogonality at every element of the space. We improve the result for Banach spaces in which the set of all smooth points forms a dense $G_{\delta}$-set by proving that a norm one linear operator that preserves orthogonality on a dense subset of the space is an isometry. We further demonstrate that if such an operator preserves orthogonality on a hyperplane not passing through the origin then it is an isometry. In the context of finite-dimensional Banach spaces, we prove that preserving orthogonality on the set all extreme points of the unit ball forces the operator to be an isometry, which substantially refines Blanco-Koldobsky-Turnšek theorem. Finally, for finite-dimensional polyhedral spaces, we establish the significance of the set of all $k$-smooth points for any possible $k,$ in the study of isometric theory.
@@
Isometry; polyhedral Banach spaces; preservation of orthogonality; norm derivatives
@@
The famous Blanco-Koldobsky-Turnšek characterization [4, 10] of isometries as norm one linear operators that preserve Birkhoff-James orthogonality at each element of the space, is one of the major achievements in the isometric theory of Banach spaces. Due to appropriate importance of this result on global orthogonality preservation, it is worth exploring local versions of this phenomenon. Recently in [21], some partial refinements of the theorem was obtained in the setting of finite-dimensional Hilbert spaces and some special finite-dimensional polyhedral spaces. The purpose of this article is to substantially refine the Blanco-Koldobsky-Turnšek characterization theorem for general Banach spaces. We now introduce the notations and terminologies to the readers which are essential to present this article.
@@
The letters $\mathbb{X},\mathbb{Y}$ denote real Banach spaces and $\mathbb{X}^{*}$ denotes the dual space of $\mathbb{X}.$ Let $B_{\mathbb{X}}$ and $S_{\mathbb{X}}$ be the unit ball and the unit sphere of $\mathbb{X}$, respectively. The convex hull of a nonempty set $D\subset\mathbb{X}$ is denoted by $\text{co}(D).$ Let us recall that an affine space is defined as a translation of a subspace of the vector space. The intersection of all affine spaces containing a nonempty set $D\subset\mathbb{X}$ is denoted by $\rm aff(D).$ By $\text{Int}_{r}~{}D,$ we denote the relative interior of a nonempty set $D\subset\mathbb{X},$ i.e., $\text{Int}_{r}~{}D=\{x\in D:\text{ there exists }\epsilon>0\text{ such that }B(x,\epsilon)\cap\rm aff(D)\subseteq D\}.$ For a nonempty convex set $D,$ let ${\rm Ext}\,D$ be the collection of all extreme points of $D.$ If for a finite-dimensional Banach space $\mathbb{X},$ ${\rm Ext}\,B_{\mathbb{X}}$ is finite then $\mathbb{X}$ is considered a polyhedral Banach space. A convex set $F\subset S_{\mathbb{X}}$ is said to be a face of $B_{\mathbb{X}}$ if for any $u,v\in S_{\mathbb{X}},~{}(1-t)u+v\in F$ implies that $u,v\in F,$ where $10<t<$ 1 . The dimension of a face $F$ is defined as the dimension of the subspace $\mathbb{V}={\rm span}\{u-v:u,v\in F\}.$ If the dimension of the face $F$ is $k,$ then $F$ is called a $k$-face of $B_{\mathbb{X}}.$ A maximal face of $B_{\mathbb{X}}$ is called a facet.
@@
For any non-zero $z\in\mathbb{X},$ we define $J(z)$ as the collection of all supporting functionals at $z,$ i.e., $J(z)=\{f\in S_{\mathbb{X}^{*}}:f(z)=\|z\|\}.$ A non-zero $z\in\mathbb{X}$ is $k$-smooth point (or order of smoothness is $k$) if ${\rm dim}\,{\rm span}\,J(z)=k,$ where $k$ is a natural number. In particular, $1$-smooth points are called smooth points. In case of an $n$-dimensional polyhedral Banach space $\mathbb{X},$ a point $z\in S_{\mathbb{X}}$ is $k$-smooth if and only if $z$ is in the relative interior of an $(n-k)$-face of $B_{\mathbb{X}}$ [24]. For more insights on smooth and $k$-smooth points in Banach spaces and their various applications, the readers can see the articles [6, 9, 11, 14, 15, 17, 22, 25]. We denote the collection of all smooth points in $\mathbb{X}$ by ${\rm Sm}\,\mathbb{X}$ and the collection of all $k$-smooth points in $\mathbb{X}$ by ${\rm k\text{-}Sm}\,\mathbb{X}.$
@@
$u\perp_{\rho_{+}}v\Longleftrightarrow\rho^{\prime}_{+}(u,v)=0,~{}u\perp_{\rho_{-}}v\Longleftrightarrow\rho^{\prime}_{-}(u,v)=0\text{ and }u\perp_{\rho}v\Longleftrightarrow\rho^{\prime}(u,v)=0,$
@@
Given $u,v\in\mathbb{X},$ $u$ is Birkhoff–James orthogonal [3, 8] to $v,$ written as $u\perp_{B}v,$ if $\|u+\lambda v\|\geq\|u\|$ for all scalars $\lambda.$ According to the Hahn-Banach theorem, for a given $u\in\mathbb{X},$ there exist enough $v\in\mathbb{X}$ such that $u\perp_{B}v.$ Let $u^{\perp_{B}}$ be the collection of all $v\in\mathbb{X}$ such that $u\perp_{B}v.$ It is straightforward to observe that the Birkhoff-James orthogonality relation is homogeneous, i.e., $u\perp_{B}v\implies\alpha u\perp_{B}\beta v,$ for all scalars $\alpha,\beta.$ In addition, in the case of an inner product space, the Birkhoff-James orthogonality coincides with the usual inner product orthogonality. For more on Birkhoff-James orthogonality, readers can consult the recently published book [13].
The letter $T$ is reserved for a bounded linear operator either from $\mathbb{X}$ to $\mathbb{Y}$ or from $\mathbb{X}$ to $\mathbb{X},$ which will be clear from the context. Consequently we write $T\in\mathbb{L}(\mathbb{X},\mathbb{Y})$ or $T\in\mathbb{L}(\mathbb{X}).$ We say that $T$ preserves $\rho_{\pm}$orthogonality and $\rho$-orthogonality if $u,v\in\mathbb{X},$ $u\perp_{\rho_{\pm}}v\implies Tu\perp_{\rho_{\pm}}Tv$ and $u\perp_{\rho}v\implies Tu\perp_{\rho}Tv,$ respectively. Similarly, we say that $T$ preserves Birkhoff-James orthogonality if for $u,v\in\mathbb{X},$ $u\perp_{B}v\Rightarrow Tu\perp_{B}Tv.$ For the local version of this property, $T$ is said to preserve Birkhoff-James orthogonality at $u\in\mathbb{X}$ if for all $v\in\mathbb{X},$ $u\perp_{B}v\Rightarrow Tu\perp_{B}Tv.$ Several recent studies [16, 19, 20, 21, 23] have demonstrated the importance of the local preservation of Birkhoff-James orthogonality in understanding the geometric properties of the underlying Banach spaces. We recall the following definition introduced in [21], which is connected with the main objectives of this article:
@@
A set $A\subset S_{\mathbb{X}}$ is called a $\mathcal{K}$-set (Koldobsky-set) if for any operator $T\in\mathbb{L}(\mathbb{X}),$ preservation of Birkhoff-James orthogonality at each point of $A$ under $T$ implies that $T$ is an isometry up to a scalar multiplication.
It is evident that identifying a $\mathcal{K}$-set $A$ for a Banach space $\mathbb{X}$, where $A\subsetneq S_{\mathbb{X}}$, will undoubtedly refine the Blanco-Koldobsky-Turnšek Theorem for the particular Banach space $\mathbb{X}$. The primary objective of this article is to explore the $\mathcal{K}$-sets in Banach spaces in which the set of all smooth points forms a dense $G_{\delta}$-set. It is clear that our approach is related to the local preservation of Birkhoff-James orthogonality. In this article, we first show that in Banach spaces in which the set of all smooth points forms a dense $G_{\delta}$-set, the set of all smooth points constitutes a $\mathcal{K}$-set. Furthermore, we establish that any dense subset of the space is a $\mathcal{K}$-set in such spaces. Next, we establish that for a Banach space whose unit ball is the convex hull of its extreme points, the set of all extreme points of the unit ball also forms a $\mathcal{K}$-set. Finally, we show that in a finite-dimensional polyhedral Banach space, the set of all $k$-smooth points is a $\mathcal{K}$-set, for any possible $k.$
@@ 2. Main Results
Let $Tx$ be smooth. Then there exists $\psi\in S_{\mathbb{Y}^{*}}$ such that $J(Tx)=\{\psi\}.$ If possible suppose that $x$ is not smooth. Then there exist two linearly independent $f$ and $g$ in $J(x)$. Therefore, there exists $y\in\mathbb{X}$ such that $y\in\ker f\setminus\ker g$. Then $\mathbb{X}={\rm span}\,\{y\}\oplus\ker g$ and so there exist $\alpha(\neq 0)\in\mathbb{R}$ and $h(\neq 0)\in\ker g$ such that $x=\alpha y+h.$ Since $T$ preserves Birkhoff-James orthogonality at $x$, it follows that $T(x^{\perp_{B}})\subset(Tx)^{\perp_{B}}$ and so $T(\ker f\cup\ker g)\subset\ker\psi.$ This implies that $\psi(Tx)=\psi(\alpha y+h)=0.$ This contradicts the fact that $J(Tx)=\{\psi\}.$ Therefore, $x$ is smooth. ∎
@@
Now, we focus on studying the $\mathcal{K}$-sets in Banach spaces in which the set of all smooth points forms a dense $G_{\delta}$-set. For this, we need the following lemma.
@@
    If ${\rm Sm}\,\mathbb{X}$ and ${\rm Sm}\,\mathbb{Y}$ are dense $G_{\delta}$-sets of $\mathbb{X}$ and $\mathbb{Y},$ respectively and $T$ is bijective then ${\rm Sm}\,\mathbb{X}\cap T^{-1}({\rm Sm}\,\mathbb{Y})$ is dense in $\mathbb{X}.$
@@
(ii) Let ${\rm Sm}\,\mathbb{X}$ and ${\rm Sm}\,\mathbb{Y}$ be dense $G_{\delta}$-sets of $\mathbb{X}$ and $\mathbb{Y},$ respectively. Then $\mathbb{Y}\setminus{\rm Sm}\,\mathbb{Y}$ is of the first category in $\mathbb{Y}.$ Suppose $T$ is bijective. Then $T$ is a homeomorphism and therefore, $T^{-1}(\mathbb{Y}\setminus{\rm Sm}\,\mathbb{Y})=\mathbb{X}\setminus T^{-1}({\rm Sm}\,\mathbb{Y})$ is of the first category in $\mathbb{X}.$ Again $\mathbb{X}\setminus{\rm Sm}\,\mathbb{X}$ is of the first category in $\mathbb{X}.$ Now,
@@
    $\rho^{\prime}_{\pm}(\alpha u,v)=\rho^{\prime}_{\pm}(u,\alpha v)=\begin{cases}\alpha\rho^{\prime}_{\pm}(u,v),&\text{ if }\alpha\geq 0\\ \alpha\rho^{\prime}_{\mp}(u,v),&\text{ if }\alpha<0.\end{cases}$
@@
We next find a sufficient condition for a bijective operator between Banach spaces in which the set of all smooth points forms a dense $G_{\delta}$-set to be a scalar multiple of an isometry.
@@
Let ${\rm Sm}\,\mathbb{X}$ and ${\rm Sm}\,\mathbb{Y}$ be dense $G_{\delta}$-sets of $\mathbb{X}$ and $\mathbb{Y},$ respectively. If a bijective operator $T\in\mathbb{L}(\mathbb{X},\mathbb{Y})$ preserves Birkhoff-James orthogonality at each point of ${\rm Sm}\,\mathbb{X}\cap T^{-1}({\rm Sm}\,\mathbb{Y})$ then $T$ is a scalar multiple of an isometry.
@@
As a consequence of the above theorem, we obtain a refinement of the Blanco-Koldobsky-Turnšek Theorem in Banach spaces in which the set of all smooth points forms a dense $G_{\delta}$-set.
@@
Let ${\rm Sm}\,\mathbb{X}$ be a dense $G_{\delta}$-set of $\mathbb{X}.$ If $T\in\mathbb{L}(\mathbb{X})$ preserves Birkhoff-James orthogonality at each point of ${\rm Sm}\,\mathbb{X}$ then $T$ is a scalar multiple of an isometry, i.e., ${\rm Sm}\,\mathbb{X}$ is a $\mathcal{K}$-set.
@@
Let ${\rm Sm}\,\mathbb{X}$ be a dense $G_{\delta}$-set of $\mathbb{X}.$ Suppose $U\subset\mathbb{X}$ is dense in $\mathbb{X}.$ Then $U$ is a $\mathcal{K}$-set.
@@
Let ${\rm Sm}\,\mathbb{X}$ be a dense $G_{\delta}$-set of $\mathbb{X}.$ Let $H\subset\mathbb{X}$ be a hyperplane that does not pass through the origin. Then $H$ is a $\mathcal{K}$-set.
@@
By combining the results established so far, we arrive at the following characterization of isometries on Banach spaces in which the set of all smooth points forms a dense $G_{\delta}$-set.
@@
Let $\mathbb{X}$ be a Banach space such that ${\rm Sm}\,\mathbb{X}$ is a dense $G_{\delta}$-set of $\mathbb{X}.$ and $T\in\mathbb{L}(\mathbb{X})$ be non-zero. Then the following conditions are equivalent:
@@
Let $\mathbb{X}$ be a Banach space such that $B_{\mathbb{X}}=\overline{\text{co}({\rm Ext}\,B_{\mathbb{X}})}$ and ${\rm Sm}\,\mathbb{X}$ is a dense $G_{\delta}$-set of $\mathbb{X}.$ If $T\in\mathbb{L}(\mathbb{X})$ preserves Birkhoff-James orthogonality on ${\rm Ext}\,B_{\mathbb{X}}$ then $T$ is a scalar multiple of an isometry, i.e., ${\rm Ext}\,B_{\mathbb{X}}$ is a $\mathcal{K}$-set.
@@
From the homogeneity of Birkhoff James orthogonality, it follows that $T$ preserves Birkhoff-James orthogonality at each point of ${\rm Sm}\,\mathbb{X}.$ Therefore, it follows from Corollary 2.8 that $T$ is an isometry multiplied by a constant. ∎
@@
    For a Banach space $\mathbb{X},$ if an operator $T\in\mathbb{L}(\mathbb{X})$ preserves $\rho$-orthogonality at each extreme point of $B_{\mathbb{X}}$ then $T$ may not be a scalar multiple of an isometry. For example, consider $\mathbb{X}=\ell_{\infty}^{2},$ then the operator $T\in\mathbb{L}(\mathbb{X})$ defined by $T(x,y)=(x+y,y-x)$ preserves $\rho$-orthogonality at each extreme point of $B_{\mathbb{X}}$ but $T$ is not a scalar multiple of an isometry. Also, from the above example it is clear that the image of an extreme point of $B_{\mathbb{X}}$ may not be an extreme point of $B_{\mathbb{X}}$ under the preservation of $\rho$-orthogonality at each extreme point of $B_{\mathbb{X}}.$
@@
Let $\|Tu\|=k.$ From Theorem 2.18, it follows that $\|Tx\|=k$ for all $x\in{\rm Int_{r}}\,F$ and $\frac{Tx}{k}\in G$ for some face $G$ of $B_{\mathbb{X}}.$ Thus, $\frac{T}{k}(F)\subset G.$ Since $F$ is a facet of $B_{\mathbb{X}},$ it follows that $F$ has $n$ linearly independent elements $x_{1},x_{2},\dots,x_{n}$ of $\mathbb{X}.$ For each $1\leq i\leq n,$ suppose that $Tx_{i}=y_{i}.$ Since $T$ is bijective, $y_{1},y_{2},\dots,y_{n}$ are linearly independent elements of $G.$ Therefore, $G$ is a facet of $B_{\mathbb{Y}}.$ Let $T_{1}=\frac{T}{k}.$ Now we show that $T_{1}({\rm Int_{r}}\,F)\subset{\rm Int_{r}}\,G.$ Let $v\in{\rm Int_{r}}\,F,$ then there exists an open set $D$ such that $v\in D\cap\rm aff(F)\subset F.$ Then
$T_{1}v\in T_{1}(D\cap\rm aff(F))=T_{1}(D)\cap T_{1}(\rm aff(F)).$
We claim that $T_{1}(\rm aff(F))=\rm aff(G).$ Since $T_{1}(F)\subset G,$ it follows that $T_{1}(\rm aff(F))\subset\rm aff(G).$ Let $y\in\rm aff(G)=\rm aff(\{y_{1},y_{2},\dots,y_{n}\}),$ then there exist $a_{i}\in\mathbb{R}$ with $\sum\limits_{1\leq i\leq n}a_{i}=1$ such that
$y=\sum\limits_{1\leq i\leq n}a_{i}y_{i}=\sum\limits_{1\leq i\leq n}a_{i}Tx_{i}=T_{1}(\sum\limits_{1\leq i\leq n}a_{i}x_{i})\subset T_{1}(\rm aff(F)).$
Thus, $T_{1}(\rm aff(F))=\rm aff(G).$ Next,
$Tv\in T_{1}(D\cap\rm aff(F))$ $=$ $T_{1}(D)\cap T_{1}(\rm aff(F))$
$=$ $T_{1}(D)\cap\rm aff(G)$
@@
Let $\mathbb{X}$ be an $n$-dimensional polyhedral Banach space and let $1\leq k\leq n.$ Let $F$ be a $k$-face of $B_{\mathbb{X}}$ and let $x\in{\rm Ext}\,F.$ Then there exists a $(k-1)$-face $G$ of $B_{\mathbb{X}}$ such that $x\in G\subset F.$

### v2

@@ Abstract.
We refine the well-known Blanco-Koldobsky-Turnšek theorem which states that a norm one linear operator defined on a Banach space is an isometry if and only if it preserves orthogonality at every element of the space. We improve the result for Banach spaces in which the set of all smooth points forms a dense $G_{\delta}$ set. We further demonstrate that if a norm one operator preserves orthogonality on a hyperplane not passing through the origin then it is an isometry. In the context of finite-dimensional Banach spaces, we prove that preserving orthogonality on the set of all extreme points of the unit ball forces the norm one operator to be an isometry, which substantially refines the Blanco-Koldobsky-Turnšek theorem. Finally, for finite-dimensional polyhedral spaces, we establish the significance of the set of all $k$-smooth points for any possible $k,$ in the study of isometric theory.
@@
Isometry; preservation of orthogonality; norm derivatives; polyhedral Banach spaces
@@
The famous Blanco-Koldobsky-Turnšek characterization [4, 10] of isometries as norm one linear operators that preserve Birkhoff-James orthogonality at each element of the space, is one of the major achievements in the isometric theory of Banach spaces. Due to appropriate importance of this result on global orthogonality preservation, it is worth exploring local versions of this phenomenon. Recently in [21], some refinements of the theorem was obtained in the setting of finite-dimensional Hilbert spaces and some special finite-dimensional polyhedral spaces. The purpose of this article is to substantially refine the Blanco-Koldobsky-Turnšek characterization theorem for general Banach spaces. We now introduce the notations and terminologies to the readers which are essential to present this article.
@@
The letters $\mathbb{X},\mathbb{Y}$ denote real Banach spaces and $\mathbb{X}^{*}$ denotes the dual space of $\mathbb{X}.$ Let $B_{\mathbb{X}}$ and $S_{\mathbb{X}}$ be the unit ball and the unit sphere of $\mathbb{X}$, respectively. The convex hull of a nonempty set $D\subset\mathbb{X}$ is denoted by $\text{co}(D).$ Let us recall that an affine space is defined as a translation of a subspace of the vector space. The intersection of all affine spaces containing a nonempty set $D\subset\mathbb{X}$ is denoted by $\text{aff}(D).$ By $\text{Int}_{r}~D,$ we denote the relative interior of a nonempty set $D\subset\mathbb{X},$ i.e., $\text{Int}_{r}~D=\{x\in D:\text{ there exists }\epsilon>0\text{ such that }B(x,\epsilon)\cap\text{aff}(D)\subseteq D\}.$ For a nonempty convex set $D,$ let ${\rm Ext}\,D$ be the collection of all extreme points of $D.$ If for a finite-dimensional Banach space $\mathbb{X},$ ${\rm Ext}\,B_{\mathbb{X}}$ is finite then $\mathbb{X}$ is considered a polyhedral Banach space. A convex subset $F$ of a convex set $G\subset\mathbb{X}$ is said to be a face of $G$ if for any $u,v\in G,~(1-t)u+v\in F$ implies that $u,v\in F,$ where $0<t<1.$ The dimension of a face $F$ is defined as the dimension of the subspace $\mathbb{V}={\rm span}\{u-v:u,v\in F\}.$ If the dimension of the face $F$ is $k,$ then $F$ is called a $k$-face of $G.$ A maximal face of $G$ is called a facet of $G$.
@@
For any non-zero $z\in\mathbb{X},$ we define $J(z)$ as the collection of all supporting functionals at $z,$ i.e., $J(z)=\{f\in S_{\mathbb{X}^{*}}:f(z)=\|z\|\}.$ A non-zero $z\in\mathbb{X}$ is $k$-smooth point (or order of smoothness is $k$) if ${\rm dim}\,{\rm span}\,J(z)=k,$ where $k$ is a natural number. In particular, $1$-smooth points are called smooth points. In case of an $n$-dimensional polyhedral Banach space $\mathbb{X},$ a point $z\in S_{\mathbb{X}}$ is $k$-smooth if and only if $z$ is in the relative interior of an (nk)(n-k)-face of $B_{\mathbb{X}}$ [24]. For more insights on smooth and $k$-smooth points in Banach spaces and their various applications, the readers can see the articles [6, 9, 11, 14, 15, 17, 22, 25]. We denote the collection of all smooth points in $\mathbb{X}$ by ${\rm Sm}\,\mathbb{X}$ and the collection of all $k$-smooth points in $\mathbb{X}$ by ${\rm k\text{-}Sm}\,\mathbb{X}.$
@@
[…]
@@
Given $v\in\mathbb{X}, uu$ is Birkhoff–James orthogonal [3, 8] to $v,$ written as $u\perp_{B}v,$ if $\|u+\lambda v\|\geq\|u\|$ for all scalars $\lambda.$ According to the Hahn-Banach theorem, for a given $u\in\mathbb{X},$ there exist enough $v\in\mathbb{X}$ such that $u\perp_{B}v.$ Let $u^{\perp_{B}}$ be the collection of all $v\in\mathbb{X}$ such that $u\perp_{B}v.$ It is straightforward to observe that the Birkhoff-James orthogonality relation is homogeneous, i.e., $u\perp_{B}v\implies\alpha u\perp_{B}\beta v,$ for all scalars $\alpha,\beta.$ In addition, in the case of an inner product space, the Birkhoff-James orthogonality coincides with the usual inner product orthogonality. For more on Birkhoff-James orthogonality, readers can consult the recently published book [13].
The letter $T$ is reserved for a bounded linear operator either from $\mathbb{X}$ to $\mathbb{Y}$ or from $\mathbb{X}$ to $\mathbb{X},$ which will be clear from the context. Consequently we write $T\in\mathbb{L}(\mathbb{X},\mathbb{Y})$ or $T\in\mathbb{L}(\mathbb{X}).$ We say that $T$ preserves $\rho_{\pm}$ -orthogonality and $\rho$-orthogonality if $u,v\in\mathbb{X},$ $u\perp_{\rho_{\pm}}v\implies Tu\perp_{\rho_{\pm}}Tv$ and $u\perp_{\rho}v\implies Tu\perp_{\rho}Tv,$ respectively. Similarly, we say that $T$ preserves Birkhoff-James orthogonality if for $u,v\in\mathbb{X},$ $u\perp_{B}v\Rightarrow Tu\perp_{B}Tv.$ For the local version of this property, $T$ is said to preserve Birkhoff-James orthogonality at $u\in\mathbb{X}$ if for all $v\in\mathbb{X},$ $u\perp_{B}v\Rightarrow Tu\perp_{B}Tv.$ Several recent studies [16, 19, 20, 21, 23] have demonstrated the importance of the local preservation of Birkhoff-James orthogonality in understanding the geometric properties of the underlying Banach spaces. We recall the following definition introduced in [21], which is connected with the main objectives of this article:
@@
A set $A\subset S_{\mathbb{X}}$ is called a $\mathcal{K}$-set (Koldobsky-set) if for any operator $T\in\mathbb{L}(\mathbb{X}),$ preservation of Birkhoff-James orthogonality at each point of $A$ under $T$ implies that $T$ is an isometry up to scalar multiplication.
It is evident that identifying a $\mathcal{K}$-set $A$ for a Banach space $\mathbb{X}$, where $A\subsetneq S_{\mathbb{X}}$, will undoubtedly refine the Blanco-Koldobsky-Turnšek theorem for the particular Banach space $\mathbb{X}$. The primary objective of this article is to explore the $\mathcal{K}$-sets in Banach spaces in which the set of all smooth points forms a dense $G_{\delta}$ set. It is clear that our approach is related to the local preservation of Birkhoff-James orthogonality. In this article, we first show that in Banach spaces in which the set of all smooth points forms a dense $G_{\delta}$ set, the set of all smooth points of the unit sphere constitutes a $\mathcal{K}$-set. Furthermore, we establish that any dense subset of the unit sphere is a $\mathcal{K}$-set in such spaces. Next, we establish that for a Banach space whose unit ball is the convex hull of its extreme points, the set of all extreme points of the unit ball also forms a $\mathcal{K}$-set. Finally, we show that in a finite-dimensional polyhedral Banach space, the set of all $k$-smooth points is a $\mathcal{K}$-set, for any possible $k.$
@@ 2. Main Results
Let $Tx$ be smooth. Then there exists $\psi\in S_{\mathbb{Y}^{*}}$ such that $J(Tx)=\{\psi\}.$ If possible suppose that $x$ is not smooth. Then there exist two linearly independent $f$ and $g$ in $J(x)$. Therefore, there exists $y\in\mathbb{X}$ such that $y\in\ker f\setminus\ker g$. Then $\mathbb{X}={\rm span}\,\{y\}\oplus\ker g$ and so there exist $\alpha(\neq 0)\in\mathbb{R}$ and $h(\neq 0)\in\ker g$ such that $x=\alpha y+h.$ Since $T$ preserves Birkhoff-James orthogonality at $x$, it follows that $T(x^{\perp_{B}})\subset(Tx)^{\perp_{B}}$ and so $T(\ker f\cup\ker g)\subset\ker\psi.$ This implies that $\psi(Tx)=\psi(\alpha Ty+Th)=0.$ This contradicts the fact that $J(Tx)=\{\psi\}.$ Therefore, $x$ is smooth. ∎
@@
Now, we focus on studying the $\mathcal{K}$-sets in Banach spaces in which the set of all smooth points forms a dense $G_{\delta}$ set. For this, we need the following lemma.
@@
    If ${\rm Sm}\,\mathbb{X}$ and ${\rm Sm}\,\mathbb{Y}$ are dense $G_{\delta}$ subsets of $\mathbb{X}$ and $\mathbb{Y},$ respectively and $T$ is bijective then ${\rm Sm}\,\mathbb{X}\cap T^{-1}({\rm Sm}\,\mathbb{Y})$ is dense in $\mathbb{X}.$
@@
(ii) Let ${\rm Sm}\,\mathbb{X}$ and ${\rm Sm}\,\mathbb{Y}$ be dense $G_{\delta}$ subsets of $\mathbb{X}$ and $\mathbb{Y},$ respectively. Then $\mathbb{Y}\setminus{\rm Sm}\,\mathbb{Y}$ is of the first category in $\mathbb{Y}.$ Suppose $T$ is bijective. Then $T$ is a homeomorphism and therefore, $T^{-1}(\mathbb{Y}\setminus{\rm Sm}\,\mathbb{Y})=\mathbb{X}\setminus T^{-1}({\rm Sm}\,\mathbb{Y})$ is of the first category in $\mathbb{X}.$ Again $\mathbb{X}\setminus{\rm Sm}\,\mathbb{X}$ is of the first category in $\mathbb{X}.$ Now,
@@
[…]
@@
We next find a sufficient condition for a bijective operator between Banach spaces in which the set of all smooth points forms a dense $G_{\delta}$ set to be a scalar multiple of an isometry.
@@
Let ${\rm Sm}\,\mathbb{X}$ and ${\rm Sm}\,\mathbb{Y}$ be dense $G_{\delta}$ subsets of $\mathbb{X}$ and $\mathbb{Y},$ respectively. If a bijective operator $T\in\mathbb{L}(\mathbb{X},\mathbb{Y})$ preserves Birkhoff-James orthogonality at each point of ${\rm Sm}\,\mathbb{X}\cap T^{-1}({\rm Sm}\,\mathbb{Y})$ then $T$ is a scalar multiple of an isometry.
@@
As a consequence of the above theorem, we obtain a refinement of the Blanco-Koldobsky-Turnšek theorem in Banach spaces in which the set of all smooth points forms a dense $G_{\delta}$ set.
@@
Let ${\rm Sm}\,\mathbb{X}$ be a dense $G_{\delta}$ subset of $\mathbb{X}.$ If $T\in\mathbb{L}(\mathbb{X})$ preserves Birkhoff-James orthogonality at each point of ${\rm Sm}\,\mathbb{X}$ then $T$ is a scalar multiple of an isometry, i.e., ${\rm Sm}\,\mathbb{X}\cap S_{\mathbb{X}}$ is a $\mathcal{K}$-set.
@@
Let ${\rm Sm}\,\mathbb{X}$ be a dense $G_{\delta}$ subset of $\mathbb{X}.$ Suppose $U\subset\mathbb{X}$ is dense in $\mathbb{X}.$ If $T\in\mathbb{L}(\mathbb{X})$ preserves Birkhoff-James orthogonality at each point of $U$ then $T$ is a scalar multiple of an isometry.
@@
Let ${\rm Sm}\,\mathbb{X}$ be a dense $G_{\delta}$ subset of $\mathbb{X}.$ Let $H\subset\mathbb{X}$ be a hyperplane that does not pass through the origin. If $T\in\mathbb{L}(\mathbb{X})$ preserves Birkhoff-James orthogonality at each point of $H$ then $T$ is a scalar multiple of an isometry.
@@
By combining the results established so far, we arrive at the following characterization of isometries on Banach spaces in which the set of all smooth points forms a dense $G_{\delta}$ set.
@@
Let $\mathbb{X}$ be a Banach space such that ${\rm Sm}\,\mathbb{X}$ is a dense $G_{\delta}$ subset of $\mathbb{X}.$ and $T\in\mathbb{L}(\mathbb{X})$ be non-zero. Then the following conditions are equivalent:
@@
Let $\mathbb{X}$ be a Banach space such that $B_{\mathbb{X}}=\overline{\text{co}({\rm Ext}\,B_{\mathbb{X}})}$ and ${\rm Sm}\,\mathbb{X}$ is a dense $G_{\delta}$ subset of $\mathbb{X}.$ If $T\in\mathbb{L}(\mathbb{X})$ preserves Birkhoff-James orthogonality on ${\rm Ext}\,B_{\mathbb{X}}$ then $T$ is a scalar multiple of an isometry, i.e., ${\rm Ext}\,B_{\mathbb{X}}$ is a $\mathcal{K}$-set.
@@
Again from the homogeneity of Birkhoff James orthogonality, it follows that $T$ preserves Birkhoff-James orthogonality at each point of ${\rm Sm}\,\mathbb{X}.$ Therefore, it follows from Corollary 2.8 that $T$ is an isometry multiplied by a constant. ∎
@@
    For a Banach space $\mathbb{X},$ if an operator $T\in\mathbb{L}(\mathbb{X})$ preserves $\rho$-orthogonality at each extreme point of $B_{\mathbb{X}}$ then $T$ may not be a scalar multiple of an isometry. For example, consider $\mathbb{X}=\ell_{\infty}^{2},$ then the operator $T\in\mathbb{L}(\mathbb{X})$ defined by T(x,y)=(x+y,yx)T(x,y)=(x+y,y-x) preserves $\rho$-orthogonality at each extreme point of $B_{\mathbb{X}}$ but $T$ is not a scalar multiple of an isometry. Also, from the example it is clear that the image of an extreme point of $B_{\mathbb{X}}$ may not be an extreme point of $B_{\mathbb{X}}$ under the preservation of $\rho$-orthogonality at each extreme point of $B_{\mathbb{X}}.$
@@
Let $\|Tu\|=k.$ From Lemma 2.16 and Theorem 2.18 (i), it follows that $\|Tx\|=k$ for all $x\in{\rm Int_{r}}\,F$ and $\frac{Tx}{k}\in G$ for some face $G$ of $B_{\mathbb{X}}.$ Thus, $\frac{T}{k}(F)\subset G.$ Since $F$ is a facet of $B_{\mathbb{X}},$ it follows that $F$ has $n$ linearly independent elements $x_{1},x_{2},\dots,x_{n}$ of $\mathbb{X}.$ For each $1\leq i\leq n,$ suppose that $Tx_{i}=y_{i}.$ Since $T$ is bijective, $y_{1},y_{2},\dots,y_{n}$ are linearly independent elements of $G.$ Therefore, $G$ is a facet of $B_{\mathbb{Y}}.$ Let $T_{1}=\frac{T}{k}.$ Now we show that $T_{1}({\rm Int_{r}}\,F)\subset{\rm Int_{r}}\,G.$ Let $v\in{\rm Int_{r}}\,F,$ then there exists an open set $D$ such that $v\in D\cap\text{aff}(F)\subset F.$ Then
$T_{1}v\in T_{1}(D\cap\text{aff}(F))=T_{1}(D)\cap T_{1}(\text{aff}(F)).$
We claim that $T_{1}(\text{aff}(F))=\text{aff}(G).$ Since $T_{1}(F)\subset G,$ it follows that $T_{1}(\text{aff}(F))\subset\text{aff}(G).$ Let $y\in\text{aff}(G)=\text{aff}(\{y_{1},y_{2},\dots,y_{n}\}),$ then there exist $a_{i}\in\mathbb{R}$ with $\sum\limits_{1\leq i\leq n}a_{i}=1$ such that
$y=\sum\limits_{1\leq i\leq n}a_{i}y_{i}=\sum\limits_{1\leq i\leq n}a_{i}Tx_{i}=T_{1}(\sum\limits_{1\leq i\leq n}a_{i}x_{i})\subset T_{1}(\text{aff}(F)).$
Thus, $T_{1}(\text{aff}(F))=\text{aff}(G).$ Next,
$Tv\in T_{1}(D\cap\text{aff}(F))$ $=$ $T_{1}(D)\cap T_{1}(\text{aff}(F))$
$=$ $T_{1}(D)\cap\text{aff}(G)$
@@
Let $\mathbb{X}$ be an $n$-dimensional polyhedral Banach space and let $1\leq k\leq n.$ Let $F$ be a $k$-face of $B_{\mathbb{X}}$ and let $x\in{\rm Ext}\,F.$ Then there exists a (k1)(k-1)-face $G$ of $B_{\mathbb{X}}$ such that $x\in G\subset F.$